Perrette Motor Company rebuilds automobile engines that have been damaged or are in need of extensive repair. The rebuilt engine has a 100,000-mile warranty and is purchased by auto shops, large motor pools in companies and governmental units, and some individual auto owners. The plant in Dayton, Ohio, specializes in the Ford V6 engine. Approximately 930 to 1,130 engines are rebuilt each month, and the costs of the plant are assigned to monthly production using weighted-average process costing. The current month began with a Work-in-Process Inventory of 320 engines, which were 50% complete for direct materials and 50% complete for conversion costs. The direct direct materials cost in beginning Work-in-Process was $124,900, while the conversion cost was $124,410. A total of 830 engines were completed and shipped out during the month, and total direct materials cost of $904,700 and conversion cost of $1,008,450 were incurred during the month. The ending Work-in-Process of 420 units was 50% complete for direct materials and 30% complete for conversion costs.

Required:

Compute the cost per equivalent unit for direct materials and conversion for the month.

                                                          Cost per equivalent unit

Direct Materials-

Conversion-
Answer #1

Equivalent units = Units completed + (Ending Work-in-Process X % completed)

Material = 830 + (420 X 50%) = 1,040 units

Conversion costs = 830 + (420 X 30%) = 956 units

Total cost:

Material = $124,900 + $904,700 = $1,029,600

Conversion cost = $124,410 + $1,008,450 = $1,132,860

Cost per equivalent unit = Total cost / Equivalent units

Material = $1,029,600 / 1,040 = $990

Conversion cost = $1,132,860 / 956 = $1,185